OpenStudy (realitywillslapyou):

Select the equations that are parallel and perpendicular to y = −3x − 1 and that pass through the point (3, 1). A. parallel: y = −3x + 6 perpendicular: y = (1/3)x + 2(2/3) B. parallel: y = −3x + 10 perpendicular: y = (1/3)x C. parallel: y = -(1/3)x perpendicular: y = −3x D. parallel: y = (1/3)x + 1 perpendicular: y = 3x − 1

OpenStudy (realitywillslapyou):

@SmokeysTheName, Right here!

OpenStudy (smokeysthename):

ohhh man, remember i said i couldn't help? but here @phi might be able to

OpenStudy (realitywillslapyou):

Oh, oops! Sorry about that.... :)

OpenStudy (smokeysthename):

thats fine (:

OpenStudy (smokeysthename):

@hero can you help with this?

hero (hero):

Pretty easy stuff @SmokeysTheName

OpenStudy (smokeysthename):

its been forever since ive done this and i can't remember /: so can you help him?

hero (hero):

Hint: m1 x m2 = -1

OpenStudy (realitywillslapyou):

Umm.... I not really sure.. :|

hero (hero):

m1 = -3 So m2 has to be a number such that when multiplied by m1, their product is -1.

OpenStudy (realitywillslapyou):

-1?

OpenStudy (realitywillslapyou):

Then 1 x -1 = -1

OpenStudy (realitywillslapyou):

Right?

hero (hero):

m1 = -3 Multiply -3 by something to get -1

OpenStudy (realitywillslapyou):

I have no clue..... :|

OpenStudy (smokeysthename):

is that even possible @hero?

hero (hero):

-3 x 1/3 = -1

hero (hero):

Yes, it is possible.

hero (hero):

So that's basically the slope of the second equation.

OpenStudy (smokeysthename):

oh yeah xD lol i dont like fractions so i didnt think of it. my bad okay ill stay out of this..

OpenStudy (realitywillslapyou):

Oh, why is math so hard for me to understand?!

hero (hero):

So now, you have the slope and the point. Use the point slope formula to figure out the value of b.

OpenStudy (realitywillslapyou):

Y U NO WORK FOR ME MATH?! (ノಠ益ಠ)ノ

OpenStudy (smokeysthename):

phi is right, khan acad vids are great to watch

OpenStudy (realitywillslapyou):

@phi, in the video he was using two points and my question only has one point so how can I solve it the same way?

OpenStudy (phi):

at the very end he does a problem like yours.

OpenStudy (realitywillslapyou):

Oh, my video sever crashed after the second equation.

OpenStudy (phi):

but this video assumes you know about slopes and equations of lines (he has videos on all of that) for you problem, you should know that when you see y = mx +b m is the slope. the number multiplying the x is the slope

OpenStudy (realitywillslapyou):

I will start it back up again, but it will take 2 min.

OpenStudy (realitywillslapyou):

I tried to solve it (I think I did it wrong) and is it C?

OpenStudy (phi):

for you problem, you should know that when you see y = mx +b m is the slope. the number multiplying the x is the slope so what is the slope of your line y = −3x − 1 ?

OpenStudy (realitywillslapyou):

-3?

OpenStudy (phi):

do you see a number in front of the x ?

OpenStudy (realitywillslapyou):

So is the answer B?

OpenStudy (phi):

yes, -3 is the slope (that is why people came up with this kind of equation, you can read off the slope easily)

OpenStudy (phi):

from the video, you learned that parallel lines have what kind of slope ?

OpenStudy (realitywillslapyou):

The same slope

OpenStudy (phi):

look at your 4 choices. which choices have the correct slope of -3 for the parallel line ?

OpenStudy (realitywillslapyou):

Well it's A or B....... I think it's B

OpenStudy (phi):

no need to guess. now for the perpendicular line. Its slope is the negative inverse (or negative reciprocal) that means "flip" the slope and multiply by -1. For example if you had slope =2 \[ \frac{2}{1} ->- \frac{1}{2}\] can you do that for -3 ?

OpenStudy (realitywillslapyou):

\[\frac{ 3 }{ 1 }?\]

OpenStudy (realitywillslapyou):

Or\[\frac{ 1 }{ 3 }?\]

OpenStudy (phi):

3/1 is 3 start with -3 or -3/1 minus it : - -3/1 is +3/1 now flip it : 1/3 you know you did it right if it is the opposite sign of -3, and it's "upside down"

OpenStudy (realitywillslapyou):

Okay

OpenStudy (phi):

the slope of the perpendicular is 1/3 look at choices A or B. can we rule out either choice because it does not have 1/3 for the slope of the perpendicular line ?

OpenStudy (realitywillslapyou):

We can rule out A, right?

OpenStudy (phi):

for choice A, what is the number in front of the x for the perpendicular line ?

OpenStudy (realitywillslapyou):

-3

OpenStudy (phi):

that is the parallel line

OpenStudy (phi):

A. parallel: y = −3x + 6 perpendicular: y = (1/3)x + 2(2/3)

OpenStudy (realitywillslapyou):

So we rule out B?

OpenStudy (phi):

B. parallel: y = −3x + 10 perpendicular: y = (1/3)x what is the slope for the perpendicular line ?

OpenStudy (realitywillslapyou):

1/3

OpenStudy (phi):

so both A and B pass. we have to use the last bit of info pass through the point (3, 1) that says that when x=3 y must = 1 test choice B, perpendicular (because it looks the easiest) y = (1/3)x replace x with 3 and do the arithmetic what do you get ?

OpenStudy (phi):

replace x with 3 that means wherever you see an x, put a 3 in its place.

OpenStudy (realitywillslapyou):

okay give me a min.

OpenStudy (realitywillslapyou):

y = (1/3)3

OpenStudy (realitywillslapyou):

??

OpenStudy (phi):

yes, \[y= \frac{1}{3} \cdot 3 \] you can think of the 3 as 3/1 or \[y= \frac{1}{3} \cdot \frac{3}{1} \] when you multiply fractions, multiply top times top and bottom times bottom you get \[y= \frac{3}{3} \] can you simplify this ?

OpenStudy (realitywillslapyou):

y = 1

OpenStudy (phi):

so (3,1) is on this line. this is the choice

OpenStudy (phi):

if we put x=3 in choice A, we will not get y=1 btw, we should check the parallel line can you do that ? parallel: y = −3x + 10 what do you get if you replace x with 3

OpenStudy (realitywillslapyou):

-9

OpenStudy (phi):

what about the +10 ?

OpenStudy (realitywillslapyou):

1

OpenStudy (phi):

y = −3x + 10 if this does not give you y=1 when x=3 then it is not the line. But it does give y=1 so choice B. Does any of this make sense ?

OpenStudy (realitywillslapyou):

Yeah....

OpenStudy (realitywillslapyou):

How many more questions can you help me with?

OpenStudy (phi):

how many do you have ?

OpenStudy (realitywillslapyou):

Three, but I'm really bad at math.

OpenStudy (phi):

did you post them yet ?

OpenStudy (realitywillslapyou):

I posted one of them in algebra